when a stone is thrown upwards with a velocity of 15 metre per second  from a cliff 200 metre high then what is its time taken to reach the ground ?
 
 

Grade:11

1 Answers

Nimish bansal
6 years ago
First we will calculate the distance travelled by stone until its velocity becomes 0 because of the retardation. V^2=u^2 + 2asAt max distamce its velocity will become 00= 15×15 + 2(-g)sTaking g=10m/ss= 45/4 m.Total distance from ground= 45/4 + 200 m.Time taken to reach maximum height isV=u + at0 = 15 -10tt= 3/2 seconds.Particle does free fall from height 45/4 +200m.S= ut + 1/2 at^2200+45/4 = 0 + 1/2×10 ×t^2t= 6.5 seconds.Total time = 6.5 + 1.5 = 8 secondsAns is t= 8 seconds